# On Bayesian Neural Networks with Dependent and Possibly Heavy-Tailed Weights

### v4

@@ 3.1 Matrices
$\mathrm{Tr}(\mathbb{A}\mathbb{B})=\mathrm{vec}(\mathbb{A}^{\top})^{\top}\mathrm{vec}(\mathbb{B}),$ (1)
it follows that ${\rm vec}:\mathbb{R}^{p\times q}\rightarrow\mathbb{R}^{pq}$ is an isometry between the Euclidean spaces $(\mathbb{R}^{p\times q},\langle\cdot,\cdot\rangle_{F})$ and $(\mathbb{R}^{pq},\langle\cdot,\cdot\rangle)$. In view of this bijection, we think of a random matrix $\mathbb{U}$ as a matrix-valued random variable whose law is induced by the random vector ${\rm vec}(\mathbb{U})$ (via ${\rm vec}^{-1})$. Accordingly, the characteristic function of the random matrix $\mathbb{U}\in\mathbb{R}^{p\times q}$ is given by
$\phi(\boldsymbol{\theta};\mathbb{U})=\mathbb{E}\left(e^{\mathbb{i}\langle\boldsymbol{\theta},\mathbb{U}\rangle_{F}}\right)=\mathbb{E}\left(e^{\mathbb{i}\mathrm{Tr}(\boldsymbol{\theta}^{\top}\mathbb{U})}\right)$
where $\mathbb{i}:=\sqrt{-1}$ and $\boldsymbol{\theta}$ ranges in $\mathbb{R}^{p\times q}$. The Kronecker (also known as tensor) product of matrices $\mathbb{A}=(a_{ij})\in\mathbb{R}^{p\times q}$ and $\mathbb{B}\in\mathbb{R}^{r\times s}$, is the $pr\times qs$ real matrix $\mathbb{A}\otimes\mathbb{B}=(a_{ij}\mathbb{B})_{1\leq i\leq p,\,1\leq j\leq q}$. If $\mathbb{A}$ and $\mathbb{B}$ are invertible matrices, then $(\mathbb{A}\otimes\mathbb{B})^{-1}=\mathbb{A}^{-1}\otimes\mathbb{B}^{-1}$. Moreover, if $\mathbb{A},\,\mathbb{B}$ and $\mathbb{C}$ are matrices such that $\mathbb{ABC}$ exists, then $\mathrm{vec}(\mathbb{ABC})=(\mathbb{C}^{\top}\otimes\mathbb{A})\mathrm{vec}(\mathbb{B})$. Therefore, with $\mathbb{Id}$ the identity matrix of appropriate dimension, choosing $\mathbb{C}=\mathbb{Id}$ and $\mathbb{A}$ invertible, identity (1) and the preceding properties yield the useful relations
$\mathrm{vec}(\mathbb{B})^{\top}(\mathbb{Id}\otimes\mathbb{A})\mathrm{vec}(\mathbb{B})=\mathrm{Tr}(\mathbb{B^{\top}AB})\,\,\text{and}\,\,\mathrm{vec}(\mathbb{B})^{\top}(\mathbb{Id}\otimes\mathbb{A})^{-1}\mathrm{vec}(\mathbb{B})=\mathrm{Tr}(\mathbb{B^{\top}A^{-1}B}).$ (2)
For later purposes, we recall that, for every symmetric positive semi-definite matrix $\mathbb{A}=(a_{rs})_{r,s=1,\ldots,d}$, there exists a unique symmetric positive semi-definite matrix $\mathbb{A}^{\sharp}=(a_{rs}^{\sharp})_{r,s=1,\ldots,d}$ (the square-root of $\mathbb{A}$) such that $\mathbb{A}^{\sharp}\mathbb{A}^{\sharp}=\mathbb{A}$, i.e., $a_{rs}=\sum_{j=1}^{d}a_{rj}^{\sharp}a_{js}^{\sharp}$, $r,s=1,\ldots,d$.
For $\mathbb{A}\in\mathbb{R}^{p\times q}$, we denote by $rk(\mathbb{A})$ the rank of $\mathbb{A}$, i.e., the number of linearly independent columns or rows within $\mathbb{A}$. We denote by 0 the null vector of $\mathbb{R}^{p}$, by ${\rm diag}_{r}(a_{1},\ldots,a_{r})$, $r\in\mathbb{N}^{*}$, a diagonal $r\times r$ real matrix with diagonal elements $a_{1},\ldots,a_{r}$, by $\mathbb{Id}_{p}$ the $p\times p$ identity matrix, and we set $1_{p}:=(1,\ldots,1)^{\top}\in\mathbb{R}^{p}$, $p\in\mathbb{N}^{*}$.
The following elementary lemma can be easily proved.
Lemma 3.1.
@@
3.2 Infinitely divisible distributions
A real-valued random variable $X$ is said to have an infinitely divisible law if, for each $n\in\mathbb{N}^{*}$, there exist independent and identically distributed $\mathbb{R}$-valued random variables $X_{n1},\ldots,X_{nn}$ such that $X\stackrel{{\scriptstyle\mathcal{L}}}{{=}}X_{n1}+\ldots+X_{nn}$.
If $X$ is a non-negative random variable it turns out that $X$ has an infinitely divisible distribution if and only if there exists a couple $(a,\rho)$, where $a\geq 0$ is a non-negative constant and $\rho$ is a Lévy measure on $(0,\infty)$, i.e., a Borel measure on $(0,\infty)$ with
@@ 3.2 Infinitely divisible distributions
such that $X$ has characteristic function of the form ${\rm e}^{\Psi(u)}$ with
$\Psi(u)=\mathbb{i}ua+\int_{(0,\infty)}(\mathrm{e}^{\mathbb{i}ux}-1)\rho(\mathrm{d}x).$
We write $X=\mathrm{ID}(a,\rho)$.
3.3 Gaussian mixtures
We denote by $\mathcal{N}_{m}(c,\mathbb{C})$ the $m$-dimensional Gaussian distribution with mean $c\in\mathbb{R}^{m}$ and covariance matrix $\mathbb{C}$. Let $\kappa$ be a random vector with values on $\mathbb{R}^{m}$ and let $\mathbb{K}$ be a positive semi-definite and symmetric random matrix with values on $\mathbb{R}^{m\times m}$. We say that a random vector $X$, defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ and with values on $\mathbb{R}^{m}$, has the Gaussian mixture distribution (under $\mathbb{P}$) with random mean $\kappa$ and random covariance matrix $\mathbb{K}$, denoted by $\mathcal{GM}_{m}^{\mathbb{P}}(\kappa,\mathbb{K})$, if $X\,|\,(\kappa,\mathbb{K})\sim\mathcal{N}_{m}(\kappa,\mathbb{K})$. Note that if $\mathbb{P}({\rm det}\mathbb{K}>0)=1$, then $X$ has density (with respect to the Lebesgue measure) $\mathbb{E}[\varphi_{(\kappa,\mathbb{K})}(\xi)]$, $\xi\in\mathbb{R}^{m}$, being $\varphi_{(c,\mathbb{C})}$ the density of $\mathcal{N}_{m}(c,\mathbb{C})$. In the definition of Gaussian mixture we explicited the dependence on the probability measure $\mathbb{P}$ (with an upper index) since in this paper we will work with Gaussian mixtures under different probability measures.
3.4 Sequential limits
For a positive integer $\ell$ let $\ell]:=\{1,\ldots,\ell\}$. We think of $\ell$] either as a set of indices or as an interval of $\mathbb{N}$, linearly ordered by the natural order. Let $f:\mathbb{N}^{\ell}\rightarrow\mathbb{R}$ be a function, $S:=\{i_{1},\ldots,i_{r}\}$, $i_{1}<i_{2}\cdots<i_{r}$ be a linearly ordered subset of $\ell$] and $\beta:S\rightarrow[\ell]$ be an injection. By $\beta\overrightarrow{\lim}f$ we mean
$\lim_{n_{\beta(i_{r})}\to\infty}\lim_{n_{\beta(i_{r-1})}\to\infty}\cdots\lim_{n_{\beta(i_{1})}\to\infty}f,$
and we say that the limit is sequential along $\beta$. Note that if $S=[\ell]$, then $\beta$ is a permutation $\pi$ of $\ell$]. In particular, if $\pi$ is the identity, then we write $\overrightarrow{\lim}$ for $\pi\overrightarrow{\lim}$. If $\pi\overrightarrow{\lim}f=\overline{f}$, then we also write $f\xhookrightarrow{\pi}\overline{f}$. If $\beta$ is the injection obtained by restricting $\pi$ over $S$, then we write $\overline{f}_{\beta}$ for the sequential limit of $f$ along $\beta$. By $(\pi\preceq j)$ we denote the injection obtained by restricting $\pi$ on $[j]$. Analogously, $(\pi\succeq j)$ denotes the injection obtained by restricting $\pi$ on $\{j,j+1,\ldots,\ell\}$. If $j=0$, then $(\pi\preceq 0)$ is the unique permutation of the empty set and we set $(\pi\preceq 0)\overrightarrow{\lim}f=f$. Clearly, if $j=\ell$, then $(\pi\preceq j)=\pi$ while $(\pi\succeq j)$ is the unique permutation of $\{\pi(\ell)\}$. Moreover, if $j\not$$\ell$, then
$\pi\overrightarrow{\lim}f=(\pi\succeq j+1)\overrightarrow{\lim}\left[(\pi\preceq j)\overrightarrow{\lim}f\right].$ (3)
The preceding definitions can be adopted other notions of limits, for instance, if $f$ a random object, we write $f\xhookrightarrow{\beta\mathcal{L}}\overline{f}_{\beta}$ to mean that $f$ converges in law to $\overline{f}_{\beta}$ sequentially along $\beta$, and we write $f\xhookrightarrow{\beta\mathbb{P}}\overline{f}_{\beta}$ to mean that $f$ converges $\mathbb{P}$-a.s. to $\overline{f}_{\beta}$ sequentially along $\beta$. When $\beta$ is the identity permutation on $\ell$], we simply omit it in the corresponding symbol.
4 The infinite-width limit under the prior
We start introducing some notation. For $\ell=1,\ldots,L+1$ and $i=1,\ldots,d$, we set
@@ 4 The infinite-width limit under the prior
and we consider the $n_{\ell}\times d$ random matrix
$\mathbb{Z}^{(\ell)}(\mathbb{x}):=(Z^{(\ell)}(x(1))\ldots Z^{(\ell)}(x(d))).$
For later purposes, we vectorialize the $n_{\ell}\times d$ random matrix $\mathbb{Z}^{(\ell)}(\mathbb{x})$ defining the $n_{\ell}d$-dimensional random vector
$Z^{(\ell)}(\mathbb{x}):={\mathrm{v}ec}((\mathbb{Z}^{(\ell)}(\mathbb{x}))^{\top}),\quad\ell=1,\ldots,L+1.$
The following Theorem 4.1 is the main result in [26] (see Theorem 16 therein). It extends to the case of dependent and possibly heavy-tailed weights the Gaussian behavior in the infinite-width limit of a neural network with a Gaussian prior (see the seminal paper [33] and the more recent contributions by [19, 25, 30]).
Theorem 4.1.
@@
$(i)$ The activation function $\sigma$ is continuous and such that
$\forall$ $z\in\mathbb{R}$, $|\sigma(z)|\leq a_{1}+a_{2}|z|^{a_{3}}$ for some positive constants $a_{1},a_{2},a_{3}>0$.
$(ii)$ $\forall$ $\ell=1,\ldots,L$, $\sum_{j=1}^{n_{\ell}}V_{n_{\ell},j}^{(\ell)}\to\mathrm{ID}(a^{(\ell)},\rho^{(\ell)})$ in distribution as $n_{\ell}\to\infty$, for some non-negative constant $a^{(\ell)}$ and Lévy measure $\rho^{(\ell)}(\cdot)$ on $(0,\infty)$.
@@
$\overrightarrow{\lim}Z^{(L+1)}(\mathbb{x})=G^{(L+1)}(\mathbb{x})\sim\mathcal{GM}_{n_{L+1}d}^{\mathbb{P}_{\rm prior}}(0,\mathbb{Id}_{n_{L+1}}\otimes\mathbb{K}^{(L+1)}(\mathbb{x}))\,\,\,\text{in distribution}.$ (4)
@@
$\mathbb{K}^{(1)}(\mathbb{x}):=(K^{(1)}(x(i),x(i^{\prime})))_{1\leq i,i^{\prime}\leq d},$ (5)
where
$K^{(1)}(x(i),x(i^{\prime})):=C_{B}+C_{W}\frac{x(i)^{\top}x(i^{\prime})}{n_{0}},$
and
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum_{j=1}^{N_{\ell-1}((0,\infty))}T_{j}^{(\ell-1)}\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))^{\top}\Biggr),$ (6)
for $\ell=2,\ldots,L+1$.
Here, $\{\zeta_{j}^{(1)}(\mathbb{x})\}_{j\geq 1}$ is a sequence of independent and identically distributed random vectors with $\zeta_{1}^{(1)}(\mathbb{x})\sim\mathcal{N}_{d}(0,\mathbb{K}^{(1)}(\mathbb{x}))$, for $\ell=3,\ldots,L+1$, given $\mathbb{K}^{(\ell-1)}(\mathbb{x})$, $\{\zeta_{j}^{(\ell-1)}(\mathbb{x})\}_{j\geq 1}$ is a sequence of independent and identically distributed random vectors with $\zeta_{1}^{(\ell-1)}(\mathbb{x})\sim\mathcal{N}_{d}(0,\mathbb{K}^{(\ell-1)}(\mathbb{x}))$, and, for $\ell=2,\ldots,L+1$, $\{T_{j}^{(\ell-1)}\}_{j=1,\ldots,N_{\ell-1}((0,\infty))}$ are the points of a Poisson process on $(0,\infty)$ with mean measure $\rho^{(\ell-1)}$, independent of the sequence $\{\zeta_{j}^{(\ell-1)}(\mathbb{x})\}_{j\geq 1}$.
Let us represent $G^{(L+1)}(\mathbb{x})$ (the limiting random vector in (4)) as
$G^{(L+1)}(\mathbb{x}):=\mathrm{vec}((\mathbb{G}^{(L+1)}(\mathbb{x}))^{\top}),$ (7)
where $\mathbb{G}^{(L+1)}(\mathbb{x})$ is the $n_{L+1}\times d$ random matrix
$\mathbb{G}^{(L+1)}(\mathbb{x}):=(G^{(L+1)}(x(1))\ldots G^{(L+1)}(x(d))),$
and
$G^{(L+1)}(x(i)):=(G_{1}^{(L+1)}(x(i)),\ldots,G_{n_{L+1}}^{(L+1)}(x(i)))^{\top},\quad i=1,\ldots,d.$
Note that, under $\mathbb{P}_{{\rm prior}}$, the rows of $\mathbb{G}^{(L+1)}(\mathbb{x})$ are pairwise dependent and uncorrelated, and identically distributed according to the Gaussian mixture law with mean 0 and random covariance matrix $\mathbb{K}^{(L+1)}(\mathbb{x})$. In Section 5 we will see that if we consider the posterior distribution with a Gaussian likelihood, then, under a suitable probability measure, the infinite-width limit of the posterior output has dependent and, in general, not identically distributed coordinates (see Theorem 5.1).
The limit in (4) not only is sequential, but it has also a prescribed order. Theorem 4.2 provides sufficient conditions which guarantee that in the sequential limit the order doesn’t matter. Its proof is quite technical and postponed in Appendix.
Theorem 4.2.
Assume $L\geq 2$ and the condition $(ii)$ of Theorem 4.1. If either
$\sigma$ is continuous and bounded (8)
@@
$\sigma$ satisfies the condition $(i)$ of Theorem 4.1 and $\rho^{(\ell)}((0,\infty))<\infty$ $\forall$ $\ell=2,\ldots,L$, (9)
@@
$\pi\overrightarrow{\lim}Z^{(L+1)}(\mathbb{x})=G^{(L+1)}(\mathbb{x})\sim\mathcal{GM}_{n_{L+1}d}^{\mathbb{P}_{\rm prior}}(0,\mathbb{Id}_{n_{L+1}}\otimes\mathbb{K}^{(L+1)}(\mathbb{x}))\,\,\,\text{in distribution}.$
Remark 4.3.
Note that, although Theorem 4.2 generalizes (under suitable assumptions) the claim (4), it does not imply that the weak convergence to a Gaussian mixture holds as $\min\{n_{1},\ldots,n_{L}\}\to\infty$.
5 The infinite-width limit under the posterior
Theorem 4.1 provides the distribution of the output of the neural network in the infinite-width limit, under the prior probability measure $\mathbb{P}_{{\rm prior}}$ defined in Section 2.2. Here we give the distribution of the output of the neural network in the infinite-width limit, under the posterior probability measure $\mathbb{P}_{{\rm posterior}}$ defined by
$\mathrm{d}\mathbb{P}_{{\rm posterior}}\propto\exp\left(-\sum_{i=1}^{d}\|Z^{(L+1)}(x(i))-y(i)\|^{2}\right)\mathrm{d}\mathbb{P}_{{\rm prior}}.$ (10)
Hereon, (when the inverse matrices exist) we will consider the $n_{L+1}d\times n_{L+1}d$ random matrix
$\boldsymbol{\Lambda}^{(L+1)}(\mathbb{x}):=\mathbb{Id}_{n_{L+1}}\otimes\mathbb{L}^{(L+1)}(\mathbb{x})^{-1},$
where $\mathbb{L}^{(L+1)}(\mathbb{x})$ is the $d\times d$ random matrix
$\mathbb{L}^{(L+1)}(\mathbb{x}):=2\mathbb{Id}_{d}+\mathbb{K}^{(L+1)}(\mathbb{x})^{-1}.$
We will also consider the $n_{L+1}d$-dimensional random vector
$\lambda^{(L+1)}(\mathbb{x},\mathbb{y}):={\rm vec}(\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y})^{\top}),$
where $\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y})$ is the $n_{L+1}\times d$ random matrix
$\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y}):=2\mathbb{y}\mathbb{L}^{(L+1)}(\mathbb{x})^{-1}.$
Hereafter, for a $p$-dimensional random element $X$ defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and with values on some metric space, we denote by $\mathbb{P}_{X}$ (or by $(\mathbb{P})_{X}$) the probability law induced by $X$ on the metric space.

### v5

@@ 3.1. Matrices
(3.1) $\mathrm{Tr}(\mathbb{A}\mathbb{B})=\mathrm{vec}(\mathbb{A}^{\top})^{\top}\mathrm{vec}(\mathbb{B}),$
it follows that ${\rm vec}:\mathbb{R}^{p\times q}\rightarrow\mathbb{R}^{pq}$ is an isometry between the Euclidean spaces $(\mathbb{R}^{p\times q},\langle\cdot,\cdot\rangle_{F})$ and $(\mathbb{R}^{pq},\langle\cdot,\cdot\rangle)$. In view of this bijection, we think of a random matrix $\mathbb{U}$ as a matrix-valued r.v. whose law is induced by the r.v. ${\rm vec}(\mathbb{U})$ (via ${\rm vec}^{-1})$. Accordingly, the characteristic function of the random matrix $\mathbb{U}\in\mathbb{R}^{p\times q}$ is given by $\phi(\boldsymbol{\theta};\mathbb{U})=\mathbb{E}\left(e^{\mathbb{i}\langle\boldsymbol{\theta},\mathbb{U}\rangle_{F}}\right)=\mathbb{E}\left(e^{\mathbb{i}\mathrm{Tr}(\boldsymbol{\theta}^{\top}\mathbb{U})}\right)$ where $\mathbb{i}:=\sqrt{-1}$ and $\boldsymbol{\theta}\in\mathbb{R}^{p\times q}$. The Kronecker (also known as tensor) product of matrices $\mathbb{A}=(a_{ij})\in\mathbb{R}^{p\times q}$ and $\mathbb{B}\in\mathbb{R}^{r\times s}$, is the $pr\times qs$ real matrix $\mathbb{A}\otimes\mathbb{B}=(a_{ij}\mathbb{B})_{1\leq i\leq p,\,1\leq j\leq q}$. If $\mathbb{A}$ and $\mathbb{B}$ are invertible matrices, then $(\mathbb{A}\otimes\mathbb{B})^{-1}=\mathbb{A}^{-1}\otimes\mathbb{B}^{-1}$. Moreover, if $\mathbb{A},\,\mathbb{B}$ and $\mathbb{C}$ are matrices such that $\mathbb{ABC}$ exists, then $\mathrm{vec}(\mathbb{ABC})=(\mathbb{C}^{\top}\otimes\mathbb{A})\mathrm{vec}(\mathbb{B})$. Therefore, with $\mathbb{Id}$ the identity matrix of appropriate dimension, choosing $\mathbb{C}=\mathbb{Id}$ and $\mathbb{A}$ invertible, identity (3.1) and the preceding properties yield the useful relations
(3.2) $\mathrm{vec}(\mathbb{B})^{\top}(\mathbb{Id}\otimes\mathbb{A})\mathrm{vec}(\mathbb{B})=\mathrm{Tr}(\mathbb{B^{\top}AB})\,\,\text{and}\,\,\mathrm{vec}(\mathbb{B})^{\top}(\mathbb{Id}\otimes\mathbb{A})^{-1}\mathrm{vec}(\mathbb{B})=\mathrm{Tr}(\mathbb{B^{\top}A^{-1}B}).$
For later purposes, we recall that, for every symmetric positive semi-definite matrix $\mathbb{A}=(a_{rs})_{r,s\in[d]}$, there exists a unique symmetric positive semi-definite matrix $\mathbb{A}^{\sharp}=(a_{rs}^{\sharp})_{r,s\in[d]}$ (the square-root of $\mathbb{A}$) such that $\mathbb{A}^{\sharp}\mathbb{A}^{\sharp}=\mathbb{A}$, i.e., $a_{rs}=\sum_{j=1}^{d}a_{rj}^{\sharp}a_{js}^{\sharp}$, $r,s\in[d]$. For $\mathbb{A}\in\mathbb{R}^{p\times q}$, we denote by $rk(\mathbb{A})$ the rank of $\mathbb{A}$, i.e., the number of linearly independent columns or rows within $\mathbb{A}$. We denote by 0 the null vector of $\mathbb{R}^{p}$, by ${\rm diag}_{r}(a_{1},\ldots,a_{r})$, $r\in\mathbb{N}^{*}$, a diagonal $r\times r$ real matrix with diagonal elements $a_{1},\ldots,a_{r}$, by $\mathbb{Id}_{p}$ the $p\times p$ identity matrix, and we set $1_{p}:=(1,\ldots,1)^{\top}\in\mathbb{R}^{p}$, $p\in\mathbb{N}^{*}$. The following lemma can be easily proved.
Lemma 1.
@@
3.2. Infinitely divisible distributions
A real-valued r.v. $X$ is said to have an infinitely divisible law if, for each $n\in\mathbb{N}^{*}$, there exist i.i.d. $\mathbb{R}$-valued r.v.’s $X_{n1},\ldots,X_{nn}$ such that $X\stackrel{{\scriptstyle\mathcal{L}}}{{=}}X_{n1}+\ldots+X_{nn}$. If $X$ is a non-negative r.v. it turns out that $X$ has an infinitely divisible distribution if and only if there exists a couple $(a,\rho)$, where $a\geq 0$ is a non-negative constant and $\rho$ is a Lévy measure on $(0,\infty)$, i.e., a Borel measure on $(0,\infty)$ with
@@ 3.2. Infinitely divisible distributions
such that $X$ has characteristic function of the form ${\rm e}^{\Psi(u)}$ with $\Psi(u)=\mathbb{i}ua+\int_{(0,\infty)}(\mathrm{e}^{\mathbb{i}ux}-1)\rho(\mathrm{d}x).$ We write $X=\mathrm{ID}(a,\rho)$.
3.3. Gaussian mixtures
We denote by $\mathcal{N}_{m}(c,\mathbb{C})$ the $m$-dimensional Gaussian distribution with mean $c\in\mathbb{R}^{m}$ and covariance matrix $\mathbb{C}$. Let $\kappa$ be a r.v. with values on $\mathbb{R}^{m}$ and let $\mathbb{K}$ be a positive semi-definite and symmetric random matrix with values on $\mathbb{R}^{m\times m}$. We say that a r.v. $X$, defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ and with values on $\mathbb{R}^{m}$, has the Gaussian mixture distribution (under $\mathbb{P}$) with random mean $\kappa$ and random covariance matrix $\mathbb{K}$, denoted by $\mathcal{GM}_{m}^{\mathbb{P}}(\kappa,\mathbb{K})$, if $X\,|\,(\kappa,\mathbb{K})\sim\mathcal{N}_{m}(\kappa,\mathbb{K})$. Note that if $\mathbb{P}({\rm det}\mathbb{K}>0)=1$, then $X$ has density (with respect to the Lebesgue measure) $\mathbb{E}[\varphi_{(\kappa,\mathbb{K})}(\xi)]$, $\xi\in\mathbb{R}^{m}$, being $\varphi_{(c,\mathbb{C})}$ the density of $\mathcal{N}_{m}(c,\mathbb{C})$. In the definition of Gaussian mixture we explicited the dependence on the probability measure $\mathbb{P}$ (with an upper index) since in this paper we will work with Gaussian mixtures under different probability measures.
4. The infinite-width limit under the prior
We start introducing some notation. For $\ell\in[L+1]$ and $i\in[d]$, we set
@@ 4. The infinite-width limit under the prior
and we consider the $n_{\ell}\times d$ random matrix $\mathbb{Z}^{(\ell)}(\mathbb{x}):=(Z^{(\ell)}(x(1))\ldots Z^{(\ell)}(x(d))).$ For later purposes, we vectorialize the $n_{\ell}\times d$ random matrix $\mathbb{Z}^{(\ell)}(\mathbb{x})$ defining the $n_{\ell}d$-dimensional random vector $Z^{(\ell)}(\mathbb{x}):={\mathrm{v}ec}((\mathbb{Z}^{(\ell)}(\mathbb{x}))^{\top}),\quad\ell\in[L+1].$ The following Theorem 1 is the main result in [26] (see Theorem 16 therein). It extends to the case of dependent and possibly heavy-tailed weights the Gaussian behavior in the infinite-width limit of a neural network with a Gaussian prior, see the seminal paper [33] and the more recent contributions [19, 25, 30].
Theorem 1.
@@
$(i)$ The activation function $\sigma$ is continuous and such that $\forall$ $z\in\mathbb{R}$, $|\sigma(z)|\leq a_{1}+a_{2}|z|^{a_{3}}$ for some positive constants $a_{1},a_{2},a_{3}>0$.
$(ii)$ $\forall$ $\ell\in[L]$, $\sum_{j=1}^{n_{\ell}}V_{n_{\ell},j}^{(\ell)}\to\mathrm{ID}(a^{(\ell)},\rho^{(\ell)})$ in distribution as $n_{\ell}\to\infty$, for some non-negative constant $a^{(\ell)}$ and Lévy measure $\rho^{(\ell)}(\cdot)$ on $(0,\infty)$.
@@
(4.1) $\lim_{n_{L}\to\infty}\ldots\lim_{n_{1}\to\infty}Z^{(L+1)}(\mathbb{x})=G^{(L+1)}(\mathbb{x})\sim\mathcal{GM}_{n_{L+1}d}^{\mathbb{P}_{\rm prior}}(0,\mathbb{Id}_{n_{L+1}}\otimes\mathbb{K}^{(L+1)}(\mathbb{x}))$ in distribution.
@@
(4.2) $\mathbb{K}^{(1)}(\mathbb{x}):=(K^{(1)}(x(i),x(i^{\prime})))_{1\leq i,i^{\prime}\leq d},$
where $K^{(1)}(x(i),x(i^{\prime})):=C_{B}+C_{W}\frac{x(i)^{\top}x(i^{\prime})}{n_{0}},$ and, for $\ell=2,\ldots,L+1$,
@@
(4.3) $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum_{j=1}^{N_{\ell-1}((0,\infty))}T_{j}^{(\ell-1)}\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))^{\top}\Biggr).$
Here, $\{\zeta_{j}^{(1)}(\mathbb{x})\}_{j\geq 1}$ is a sequence of i.i.d. r.v.’s with $\zeta_{1}^{(1)}(\mathbb{x})\sim\mathcal{N}_{d}(0,\mathbb{K}^{(1)}(\mathbb{x}))$, for $\ell=3,\ldots,L+1$, given $\mathbb{K}^{(\ell-1)}(\mathbb{x})$, $\{\zeta_{j}^{(\ell-1)}(\mathbb{x})\}_{j\geq 1}$ is a sequence of i.i.d. r.v.’s with $\zeta_{1}^{(\ell-1)}(\mathbb{x})\sim\mathcal{N}_{d}(0,\mathbb{K}^{(\ell-1)}(\mathbb{x}))$, and, for $\ell=2,\ldots,L+1$, $\{T_{j}^{(\ell-1)}\}_{j=1,\ldots,N_{\ell-1}((0,\infty))}$ are the points of a Poisson process on $(0,\infty)$ with mean measure $\rho^{(\ell-1)}$, independent of the sequence $\{\zeta_{j}^{(\ell-1)}(\mathbb{x})\}_{j\geq 1}$.
Let us represent $G^{(L+1)}(\mathbb{x})$ (the limiting random vector in (4.1)) as
(4.4) $G^{(L+1)}(\mathbb{x}):=\mathrm{vec}((\mathbb{G}^{(L+1)}(\mathbb{x}))^{\top}),$
where $\mathbb{G}^{(L+1)}(\mathbb{x})$ is the $n_{L+1}\times d$ random matrix $\mathbb{G}^{(L+1)}(\mathbb{x}):=(G^{(L+1)}(x(1))\ldots G^{(L+1)}(x(d))),$ and $G^{(L+1)}(x(i)):=(G_{1}^{(L+1)}(x(i)),\ldots,G_{n_{L+1}}^{(L+1)}(x(i)))^{\top},\quad i\in[d].$ Note that, under $\mathbb{P}_{{\rm prior}}$, the rows of $\mathbb{G}^{(L+1)}(\mathbb{x})$ are pairwise dependent and uncorrelated, and identically distributed according to the Gaussian mixture law with mean 0 and random covariance matrix $\mathbb{K}^{(L+1)}(\mathbb{x})$. In Section 5 we will see that if we consider the posterior distribution with a Gaussian likelihood, then, under a suitable probability measure, the infinite-width limit of the posterior output has dependent and, in general, not identically distributed coordinates (see Theorem 3).
The limit in (4.1) not only is sequential, but it has also a prescribed order. Theorem 2 provides sufficient conditions which guarantee that in the sequential limit the order doesn’t matter. Its proof is quite technical and postponed in Appendix.
Theorem 2.
Assume $L\geq 2$ and the condition $(ii)$ of Theorem 1. If either
(4.5) $\sigma$ is continuous and bounded
@@
(4.6) $\sigma$ satisfies the condition $(i)$ of Theorem 1 and $\rho^{(\ell)}((0,\infty))<\infty$ $\forall$ $\ell=2,\ldots,L$,
@@
$\pi\lim Z^{(L+1)}(\mathbb{x})=G^{(L+1)}(\mathbb{x})\sim\mathcal{GM}_{n_{L+1}d}^{\mathbb{P}_{\rm prior}}(0,\mathbb{Id}_{n_{L+1}}\otimes\mathbb{K}^{(L+1)}(\mathbb{x}))$ in distribution
i.e., in (4.1) one can replace $\lim_{n_{L}\to\infty}\lim_{n_{L-1}\to\infty}\ldots\lim_{n_{1}\to\infty}$ with
$\pi\lim:=\lim_{n_{\pi(L)}\to\infty}\lim_{n_{\pi(L-1)}\to\infty}\ldots\lim_{n_{\pi(1)}\to\infty}.$
5. The infinite-width limit under the posterior
Theorem 1 provides the distribution of the output of the neural network in the infinite-width limit, under the prior probability measure $\mathbb{P}_{{\rm prior}}$ defined in Section 2.2. Here we give the distribution of the output of the neural network in the infinite-width limit, under the posterior probability measure $\mathbb{P}_{{\rm posterior}}$ defined by
(5.1) $\mathrm{d}\mathbb{P}_{{\rm posterior}}\propto\exp\left(-\sum_{i=1}^{d}\|Z^{(L+1)}(x(i))-y(i)\|^{2}\right)\mathrm{d}\mathbb{P}_{{\rm prior}}.$
Hereon, (when the inverse matrices exist) we will consider the $n_{L+1}d\times n_{L+1}d$ random matrix $\boldsymbol{\Lambda}^{(L+1)}(\mathbb{x}):=\mathbb{Id}_{n_{L+1}}\otimes\mathbb{L}^{(L+1)}(\mathbb{x})^{-1},$ where $\mathbb{L}^{(L+1)}(\mathbb{x})$ is the $d\times d$ random matrix $\mathbb{L}^{(L+1)}(\mathbb{x}):=2\mathbb{Id}_{d}+\mathbb{K}^{(L+1)}(\mathbb{x})^{-1}.$ We will also consider the $n_{L+1}d$-dimensional r.v. $\lambda^{(L+1)}(\mathbb{x},\mathbb{y}):={\rm vec}(\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y})^{\top}),$ where $\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y})$ is the $n_{L+1}\times d$ random matrix $\boldsymbol{\lambda}^{(L+1)}(\mathbb{x},\mathbb{y}):=2\mathbb{y}\mathbb{L}^{(L+1)}(\mathbb{x})^{-1}.$ Hereafter, for a $p$-dimensional r.v. $X$ defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and with values on some metric space, we denote by $\mathbb{P}_{X}$ (or sometimes by $(\mathbb{P})_{X}$) the probability law induced by $X$ on the metric space.